# Controlling noisy herds

### v2

@@ Discussion
We demonstrate that for a specific class of control problems – herding of indecisive collectives – temporality can significantly reduce control energy, even without prior knowledge of topology changes. Herding differs from standard control problems in complex networks in two fundamental ways. First, while traditional controllability in the context of complex networks involves steering the system from any initial state to any desired state within the state space [8, 53], herding focuses on guiding all agents (nodes) to a fixed consensus state along a predefined trajectory. Second, unlike standard control, where the controller can choose which nodes receive input [8], in herding, the controller has no choice over which node receives the control input.
@@
While we presented a simplified model to explore the effects of sheep indecisiveness in sheep-dog-handler interactions, the real-world dynamics of this system are far more intricate. Shepherd dogs can instinctively predict sheep movements, but expertly trained dogs uniquely integrate instinct with handler commands to achieve precise coordination. In successful teams, the handler and dog operate cohesively, eliminating the need for constant monitoring. Instead, they function as a unified entity, sharing cognitive resources to analyze and anticipate the sheep’s behavior in real time [23, 57, 24]. Systematically studying these sophisticated interactions – spanning verbal, physical, and visual modalities – could inspire advances in swarms robotics. Embedding distributed cognitive mechanisms across diverse agents, including drones, quadruped robots, and human or AI operators, could enable more adaptive, robust, and noise-tolerant systems capable of handling internal and external uncertainties effectively [58, 59].

### v3

@@ Discussion
We demonstrate that for a specific class of control problems–herding–temporality can significantly reduce control energy, even without prior knowledge of topology changes. While traditional controllability in the context of complex networks involves the ability of the controller in steering the system from any initial state to any desired state within the state space [8, 62], herdability focuses on guiding all agents (nodes) to a fixed consensus state along a predefined trajectory [54].
@@
While we presented a simplified model to explore the effects of sheep indecisiveness in sheep-dog-handler interactions, the real-world dynamics of this system are far more intricate. Shepherd dogs can instinctively predict sheep movements, but expertly trained dogs uniquely integrate instinct with handler commands to achieve precise coordination. In successful teams, the handler and dog operate cohesively, eliminating the need for constant monitoring. Instead, they function as a unified entity, sharing cognitive resources to analyze and anticipate the sheep’s behavior in real time [34, 66, 35]. Systematically studying these interactions, spanning verbal, physical, and visual modalities, could reveal more rich complexities hidden in these multi-species control dynamics, and offering insights into principles of decentralized and stochastic collective control.